### v1

# Multi-Field Adaptive Retrieval

Abstract
Document retrieval for tasks such as search and retrieval-augmented generation typically involves datasets that are unstructured: free-form text without explicit internal structure in each document. However, documents can have a structured form, consisting of fields such as an article title, message body, or HTML header. To address this gap, we introduce Multi-Field Adaptive Retrieval (mFAR), a flexible framework that accommodates any number of and any type of document indices on structured data. Our framework consists of two main steps: (1) the decomposition of an existing document into fields, each indexed independently through dense and lexical methods, and (2) learning a model which adaptively predicts the importance of a field by conditioning on the document query, allowing on-the-fly weighting of the most likely field(s). We find that our approach allows for the optimized use of dense versus lexical representations across field types, significantly improves in document ranking over a number of existing retrievers, and achieves state-of-the-art performance for multi-field structured data.
1 Introduction
The task of document retrieval has many traditional applications, like web search or question answering, but there has also been renewed spotlight on it as part of LLM workflows, like retrieval-augmented generation (RAG). One area of study is focused on increasing the complexity and naturalness of queries (Yang et al., 2018; Qi et al., 2019; Jeong et al., 2024; Lin et al., 2023). Another less studied area considers increased complexity of the documents (Jiang et al., 2024; Wu et al., 2024b). This represents a challenge forward from prior datasets for retrieval, like MS MARCO (Nguyen et al., 2016), which contain chunks of text that are highly related to the query. As a result, dense retrievers achieve success by embedding text into a vector representations (Karpukhin et al., 2020; Ni et al., 2022; Izacard et al., 2022) or searching over the documents via lexical match (Robertson et al., 1994). Relatedly, there are also prior methods (Gao et al., 2021; Chen et al., 2022) concerning the benefits of a hybrid representation, but these hybrid options are not mainstream solutions. We revisit both hybrid models and methods for retrieval of more complex documents.
Our motivation for this direction derives from two observations: 1) documents do have structure: fields like titles, timestamps, headers, authors, etc. and queries can refer directly to this structure; and 2) a different scoring method may be beneficial for each of these fields, as not every field is necessary to answer each query. More specifically, our goal is to investigate retrieval on structured data. Existing work on retrieval for structured data with dense representations focus on directly embedding structured knowledge into the model through pretraining approaches (Li et al., 2023; Su et al., 2023), but we would like a method which can more flexibly combine existing pretrained models and scorers. Similarly, there has been prior interest in multi-field retrieval, although those focused on retrieval with lexical or sparse features or early neural models (Robertson et al., 1994; Zaragoza et al., 2004; Zamani et al., 2018).
In this work, we demonstrate how multi-field documents can be represented through paired views on a per-field basis, with a learned mechanism that maps queries to weighted combinations of these views. Our method, Multi-Field Adaptive Retrieval (mFAR), is a retrieval approach that can accommodate any number of fields and any number of scorers (such as one lexical and one vector-based) for each field. Additionally, we introduce a lightweight component that adaptively weights the most likely fields, conditioned on the query. This allows us to overspecify the fields initially and let the model decide which scorer to use to score each field. mFAR obtains significant performance gains over existing state-of-the-art baselines. Unlike prior work, our simple approach does not require pretraining and offers some controllability at test-time. Concretely, our contributions are:
  1. 1.
    We introduce a novel framework for document retrieval, mFAR, that is aimed at structured data with any number of fields. Notably, mFAR is able to mix lexical and vector-based scorers between the query and the document’s fields.
  2. 2.
    We find that a hybrid mixture of scorers performs better than using dense or lexical-based scorers alone; we also find that encoding documents with our multi-field approach can result in better performance than encoding the entire document as a whole. As a result, mFAR achieves state-of-the-art performance on STaRK, a dataset for structured document retrieval.
  3. 3.
    We introduce an adaptive weighting technique that conditions on the query, weighting more the fields most related to the query and weighting less the fields that are less important.
  4. 4.
    Finally, we analyze the performance of our model trained from our framework; we control the availability of scorers at test-time in an ablation study to measure the importance of the individual fields in the corpus
Dataset Example Query Example Document
MS MARCO aleve maximum dose You should take one tablet every 8 to 10 hours until symptoms abate, …
BioASQ What is Piebaldism? Piebaldism is a rare autosomal dominant disorder of melanocyte development characterized by a congenital white forelock and multiple …
STaRK-Amazon Looking for a chess strategy guide from The House of Staunton that offers tactics against Old Indian and Modern defenses. Any recommendations? Title: Beating the King’s Indian and Benoni Defense with 5. Bd3
Brand: The House of Staunton
Description: … This book also tells you how to play against the Old Indian and Modern defenses.
Reviews: [{reviewerID: 1234, text:…}, {reviewerID: 1235, text:…}, …]
STaRK-MAG Does any research from the Indian Maritime University touch upon Fe II energy level transitions within the scope of Configuration Interaction? Title: Radiative transition rates for the forbidden lines in Fe II
Abstract: We report electric quadrupole and magnetic dipole transitions among the levels belonging to 3d 6 4s, 3d 7 and 3d 5 4s 2 configurations of Fe II in a large scale configuration interaction (CI) calculation. …
Authors: N.C. Deb, A Hibbert (Indian Maritime University)
STaRK-Prime What drugs target the CYP3A4 enzyme and are used to treat strongyloidiasis? Name: Ivermectin
Entity Type: drug
Details: {Description: Ivermectin is a broad-spectrum anti-parasite medication. It was first marketed under…, Half Life: 16 hours}
Target: gene/protein
Indication: For the treatment of intestinal strongyloidiasis due to …
Category: [Cytochrome P-450 CYP3A Inducers, Lactones, …]
Figure 1: Traditional documents for retrieval (top), like in MS MARCO (Nguyen et al., 2016) and BioASQ (Nentidis et al., 2023), are unstructured: free-form text that tends to directly answer the queries. Documents in the STaRK datasets (bottom) (Wu et al., 2024b), are structured: each contains multiple fields. The queries require information from some of these fields, so it is important to both aggregate evidence across multiple fields while ignoring irrelevant ones.
2 Multi-field Retrieval
While structured documents is a broad term in general, in this work, we focus on documents that can be decomposed into fields, where each field has a name and a value. As an example in Figure 1, for the STaRK-Prime document, Entity Type would be a field name and its value would be “drug.” The values themselves can have additional nested structure, like Category has a list of terms as its value. Note that this formulation of structured multi-field document is broadly applicable, as it not only includes objects like knowledge base entries, but also free-form text (chat messages, emails) along with their associated metadata (timestamps, sender, etc).
Formally, we consider a corpus of documents $\mathcal{C}$ = {$d_{1}$, $d_{2}$, …, $d_{n}$} and a set of associated fields $\mathcal{F}$ = {$f_{1}$, $f_{2}$, …, $f_{m}$} that make up each document $d$, i.e., $d=\{f:x_{f}\mid f\in\mathcal{F}\}$, where $x_{f}$ is the value for that field. Then, given a natural-language query $q$, the goal is a scoring function $s(q,d)$ that can be used to rank the documents in $\mathcal{C}$ such that the relevant documents to $q$ are at the top (or within the top-$k$). $q$ may query about values from any subset of fields, either lexically or semantically.
Figure 2: Document D and query Q are examples from the STaRK-MAG dataset. Parts of the query (highlighted) correspond with specific fields from D. Traditional retrievers (A) would score the entire document against the query (e.g. through vector similarity). In (B), our method, mFAR, first decomposes D into fields and scores each field separately against the query using both lexical- and vector-based scorers. This yields a pair of field-specific similarity scores, which are combined using our adaptive query conditioning approach to produce a document-level similarity score.
2.1 Standard Retriever and Contrastive Loss
Traditionally, $d$ is indexed in its entirety. The retriever can employ either a lexical (Robertson et al., 1994) or dense (embedding-based) (Lee et al., 2019; Karpukhin et al., 2020) scorer. A lexical scorer like BM25 (Robertson et al., 1994) directly computes $s(q,d)$ based on term frequencies. For a dense scorer, document and query encoders are used to embed $d$ and $q$, and a simple similarity function, in our case an unnormalized dot product, is used to compute $s(q,d)$
The document and query encoders can be finetuned by using a contrastive loss (Izacard et al., 2022), which aims to separate a positive (relevant) document $d_{i}^{+}$ against $k$ negative (irrelevant) documents $d_{i}^{-}$ for a given query $q$. In prior work, a shared encoder for the documents and queries is trained using this loss, and a temperature $\tau$ is used for training stability:
$\mathcal{L}_{c}=-\log\frac{e^{s(q_{i},d_{i}^{+})/\tau}}{e^{s(q_{i},d_{i}^{+})/\tau}+\sum_{d_{i}^{-}}e^{s(q_{i},d_{i}^{-})/\tau}}$ (1)
$\mathcal{L}_{c}$ is the basic contrastive loss which maximizes $P(d_{i}^{+}\mid q_{i})$. Following Henderson et al. (2017) and Chen et al. (2020a), we employ in-batch negatives to efficiently sample those negative documents by treating the other positive documents $d_{j}^{+}$ where $j\neq i$ and $1\leq j\leq k$ as part of the negative set, $d_{i}^{-}$. Furthermore, following prior work (Yang et al., 2019; Ni et al., 2022), we can include a bi-directional loss for $P(q_{i}\mid d_{i}^{+})$. Here, for a given positive document $d_{j}^{+}$, $q_{j}$ is the positive query and the other queries $q_{i},i\neq j$ become negative queries:
$\mathcal{L}_{r}=-\log\frac{e^{s(q_{i},d_{i}^{+})/\tau}}{e^{s(q_{i},d_{i}^{+})/\tau}+\sum_{q_{j},j\neq i}e^{s(q_{j},d_{i}^{+})/\tau}}$ (2)
The final loss for the (shared) encoder is $\mathcal{L}=\mathcal{L}_{c}+\mathcal{L}_{r}$.
2.2 mFAR: A Multi-Field Adaptive Retriever
Because structured documents can be decomposed into individual fields ($d=\{x_{f}\}_{f\in\mathcal{F}}$), we can score the query against each field separately. This score could be computed via lexical or dense (vector-based) methods. This motivates a modification to the standard setup above, where $s(q,d)$ can instead be determined as a weighted combination of field-wise scores and scoring methods,
$s(q,d)=\sum_{f\in\mathcal{F}}\sum_{m\in\mathcal{M}}w_{f}^{m}s_{f}^{m}(q,x_{f}).$ (3)
Here, $s_{f}^{m}(q,x_{f})$ is the score between $q$ and field $f$ of $d$ using scoring method $m$, and $\mathcal{M}$ is the set of scoring methods. For a hybrid model, =$\mathcal{M}=\{\text{lexical},\text{dense}\}$. $w_{f}^{m}$ is a weight, possibly learned, that is associated with field $f$ and scoring method $m$.
Adaptive field selection.
As presented, our method uses weights, $w_{f}^{m}$, that are learned for each field and scorer. This is useful in practice, as not every field in the corpus is useful or even asked about, like ISBN numbers or internal identifiers. Additionally, queries usually ask about information contained in a small number of fields and these fields change depending on the query.
This motivates conditioning the value of $w_{f}^{m}$ also on $q$ so that the weights can adapt to the given query by using the query text to determine the most important fields. We define an adaptation function $G$ and let $w_{f}^{m}=G(q,f,m)$. Now, the query-conditioned, or adaptive, sum is:
$s(q,d)=\sum_{f\in\mathcal{F}}\sum_{m\in\mathcal{M}}G(q,f,m)\cdot s_{f}^{m}(q,x_{f}).$ (4)
This can be implemented by learning embeddings for each field and scoring method, $\mathbf{a}_{f}^{m}$, and so $G(q,f,m)={\mathbf{a}_{f}^{m}}^{\top}\mathbf{q}$. We find that learning is more stable if we insert a nonlinearity: $G(q,f,m)=\text{softmax}(\{{\mathbf{a}_{f}^{m}}^{\top}\mathbf{q}\})$, and so this is the final version we use in mFAR.
Multiple scorers and normalization.
An objective of ours is to seamlessly incorporate scorers using different methods (lexical and dense). However, the distribution of possible scores per scorer can be on different scales. While $G$ can technically learn to normalize, we want $G$ to focus on query-conditioning. Instead, we experiment with using batch normalization (Ioffe & Szegedy, 2015) per field, that whitens the scores and learns new scalars $\gamma_{f}^{m}$ and $\beta_{f}^{m}$ for each field and scorer. Because these scores are ultimately used in the softmax of the contrastive loss, $\gamma_{f}^{m}$ acts like a bias term which modulates the importance of each score and $\beta_{f}^{m}$ has no effect.
Note that the score whitening process is not obviously beneficial or necessary, especially if the scorers do already share a similar distribution (i.e. if we only use dense scorers). We leave the inclusion of normalization as a hyperparameter included in our search.
Inference
At test time, the goal is to rank documents by $s(q,d)$ such that the relevant (gold) documents are highest. Because it can be slow to compute $|\mathcal{F}||\mathcal{M}||\mathcal{C}|$ scores for the whole corpus, we use an approximation. We first determine a top-$k$ shortlist, $\mathcal{C}_{f}^{m}$, of documents for each field and scorer and only compute the full scores for all $\bigcup_{f\in\mathcal{F},m\in\mathcal{M}}\mathcal{C}_{f}^{m}$, which results in the final ranking. Note this inexact approximation of the top-$k$ document is distinct from traditional late-stage re-ranking methods that rescore the query with each document, which is not the focus of this work.
3 Experiments
Our experiments are motivated by the following hypotheses:
  1. 1.
    Taking advantage of the multi-field document structure will lead to better accuracy than treating the document in its entirely, as a single field.
  2. 2.
    Hybrid (a combination of lexical and dense) approaches to modeling will perform better than using only one or other.
Table 1: The corpus size, number of fields, and queries (by split) for each of the STaRK datasets. For field information, refer to Table 6 in the Appendix.
Dataset Domain Num. Documents Num. Fields Train Dev. Test.
Amazon products, product reviews 950K 8 6K 1.5K 1.5K
MAG science papers, authors 700K 5 8K 2.6K 2.6K
Prime biomedical entities 130K 22 6.1K 2.2K 2.8K
3.1 Data
We use STaRK (Wu et al., 2024b), a collection of three retrieval datasets in the domains of product reviews (Amazon), academic articles (MAG), and biomedical knowledge (Prime), each derived from knowledge graphs. Amazon contains queries and documents from Amazon Product Reviews (He & McAuley, 2016) and Amazon Question and Answer Data (McAuley et al., 2015). MAG contains queries and documents about academic papers, sourced from the Microsoft Academic Graph (Wang et al., 2020), obgn-MAG, and obgn-papers100M (Hu et al., 2020). Prime contains queries and documents regarding biomedicine from PrimeKG (Chandak et al., 2022). These datasets are formulated as knowledge graphs in STaRK and are accompanied by complex queries.
In the retrieval baselines (Wu et al., 2024b), node information is linearized into documents that can be encoded and retrieved via dense methods. We likewise treat each node as a document. In our work, we preserve each node property or relation as a distinct field for our multi-field models or likewise reformat to a human-readable document for our single-field models. Compared to Amazon and MAG, we notice that Prime contains a higher number of relation types, i.e. relatively more fields in Prime are derived from knowledge-graph relations than either Amazon or MAG, where document content is derived from a node’s properties. The corpus sizes and number of fields are listed in Table 1. For more details on dataset preprocessing and the exact the fields used for each dataset, see Appendix A.
We use trec_eval111 https://github.com/usnistgov/trec_eval. for evaluation and follow Wu et al. (2024b) by reporting Hit@$1$, Hit@$5$, Recall@$20$, and mean reciprocal rank (MRR). [^2] We omit Hit@$5$ in the main paper due to space (it is reported in Appendix C).
3.2 Baselines and Prior Work
We compare primarily to prior work on STaRK, which is a set of baselines established by Wu et al. (2024b) and more recent work by Wu et al. (2024a). Specifically, they include two vector similarity search methods that use OpenAI’s text-embedding-ada-002 model, ada-002 and multi-ada-002. Notably, the latter is also a multi-vector approach, although it only uses two vectors per document: one to capture node properties and one for relational information. We also include their two LLM-based re-ranking baselines on top of ada-002.333Note their re-ranking baselines are evaluated on a random 10% subset of the data due to cost. Although our work does not perform re-ranking, we add these results to show the superiority of finetuning smaller retrievers over using generalist large language models for reranking.
More recently, AvaTaR (Wu et al., 2024a) is an agent-based method which iteratively generates prompts to improve reasoning and scoring of documents. While not comparable with our work, which does not focus on agents nor use models as large, it is the state-of-the-art method for STaRK.
Finally, we use a state-of-the-art the pretrained retrieval encoder, Contriever finetuned on MS MARCO (Izacard et al., 2022, facebook/contriever-msmarco) as a baseline for our dense scorer, which we subsequently continue finetuning on STaRK. We use BM25 (Robertson et al., 2004; Lù, 2024) as a lexical baseline. These use the single-field formatting described Section 3.1.
3.3 Experimental Setup
mFAR affords a combination of lexical and dense scorers across experiments. Like our baselines, we use BM25 as our lexical scorer and Contriever as our dense scorer. Because of potential differences across datasets, we initially consider four configurations that take advantage of mFAR’s ability to accommodate multiple fields or scorers: mFAR Dense uses all fields and the dense scorer, mFAR Lexical uses all fields and the lexical scorer, mFAR All uses all fields and both scorers, and mFAR 2 uses both scorers but the single-field (Sec. 3.1) document representation. Based on the initial results, we additionally consider mFAR 1+n, which consists of a single-field lexical scorer and multi-field dense scorer. This results in models that use $|\mathcal{F}|$, $|\mathcal{F}|$, $2|\mathcal{F}|$, $2$, and $|\mathcal{F}|+1$ scorers respectively. For each model and each dataset, we run a grid search over learning rates and whether to normalize and select the best model based on the development set.
Because Contriever is based on a 512-token window, we prioritize maximizing this window size for each field, which ultimately reduces the batch size we can select for each dataset. We use effective batch sizes of 96 for Amazon and Prime, and 192 for MAG, and train on 8xA100 GPUs. We use patience-based early stopping based on the validation loss. More details on the exact hyperparameters for each run are in Appendix B.
Table 2: Comparing our method (mFAR) against baselines and state-of-the-art methods on the STaRK test sets. ada-002 and multi-ada-002 are based on vector similarity; +{Claude3, GPT4} further adds an LLM reranking step on top of ada-002. AvaTaR is an agent-based iterative framework. Contriever-FT is a finetuned Contriever model, which is also the encoder finetuned in mFAR. mFAR is superior against prior methods and datasets, and earns a substantial margin on average across the benchmark. * In Wu et al. (2024b), these are only run on a random 10% subset.
Amazon MAG Prime Average
Model H@1 R@20 MRR H@1 R@20 MRR H@1 R@20 MRR H@1 R@20 MRR
ada-002 0.392 0.533 0.542 0.291 0.484 0.386 0.126 0.360 0.214 0.270 0.459 0.381
multi-ada-002 0.401 0.551 0.516 0.259 0.508 0.369 0.151 0.381 0.235 0.270 0.480 0.373
Claude3 reranker* 0.455 0.538 0.559 0.365 0.484 0.442 0.178 0.356 0.263 0.333 0.459 0.421
GPT4 reranker* 0.448 0.554 0.557 0.409 0.486 0.490 0.183 0.341 0.266 0.347 0.460 0.465
AvaTaR (agent) 0.499 0.606 0.587 0.444 0.506 0.512 0.184 0.393 0.267 0.376 0.502 0.455
BM25 0.483 0.584 0.589 0.471 0.689 0.572 0.167 0.410 0.255 0.374 0.561 0.462
Contriever-FT 0.383 0.530 0.497 0.371 0.578 0.475 0.325 0.600 0.427 0.360 0.569 0.467
mFAR Lexical 0.332 0.491 0.443 0.429 0.657 0.522 0.257 0.500 0.347 0.339 0.549 0.437
mFAR Dense 0.390 0.555 0.512 0.467 0.669 0.564 0.375 0.698 0.485 0.411 0.641 0.520
mFAR 2 0.574 0.663 0.681 0.503 0.721 0.603 0.227 0.495 0.327 0.435 0.626 0.537
mFAR All 0.412 0.585 0.542 0.490 0.717 0.582 0.409 0.683 0.512 0.437 0.662 0.545
mFAR 1+n 0.565 0.659 0.674 0.511 0.748 0.611 0.359 0.650 0.469 0.478 0.686 0.585
4 Results
We report the results from our mFAR models in Table 2, compared against against prior methods and baselines. Our best models—both make use of both scorers—perform significantly better than prior work and baselines: mFAR 2 on Amazon and MAG, and mFAR All on Prime and STaRK average. This includes surpassing re-ranking based methods and the strongest agentic method, AvaTaR. mFAR All performs particularly well on Prime (+20% for H@$1$). Comparatively, all models based on ada-002 have extended context windows of 2K tokens, but mFAR, with an encoder that has a much smaller context window size (512), still performs significantly better. Furthermore, our gains cannot be only attributed to finetuning or full reliance on lexical scorers since the mFAR models perform better against the already competitive BM25 and finetuned Contriever baselines.
We find that the adoption of a hybrid approach benefits recall, which we can attribute to successful integration of BM25’s scores. Individually, BM25 already achieves higher R@20 than most vector-based methods. The mFAR models retain and further improve on that performance. Recall is especially salient for tasks such as RAG where collecting documents in the top-$k$ are more important than surfacing the correct result at the top.
Revisiting our hypotheses from Section 3, we can compare the various configurations of mFAR. Noting that BM25 is akin to a single-field, lexical baseline and Contriever-FT is a single-field, dense baseline, we can observe the following:
Multi-field vs. Single-field.
A side-by-side comparison of the single-field models against their multi-field counterparts shows mixed results. If we only consider dense scorers, mFAR Dense produces the better results than Contriever-FT across all datasets. To our knowledge, this is the first positive evidence in favor of multi-field methods in dense retrieval. For mFAR Lexical, in both Amazon and MAG, the BM25 baseline performs especially well, and we do not see consistent improvements. This specific phenomenon has been previously noted by Robertson et al. (2004), who describe a modification that should be made to the BM25 algorithm itself to properly aggregate multi-field information. Specifically in BM25, the scores are length normalized. For some terms, like institution, repetition does not imply a stronger match, and so treating the institution field separately (and predicting high weights for it) could lead to high scores for negative documents. A multi-field sparse representation, then, may not always be the best solution, depending on the dataset.
Hybrid is best.
Across both multi-field (mFAR All vs. mFAR Dense or mFAR Lexical) and single-field models (mFAR 2 vs. BM25 or Contriever-FT), and across almost every dataset, there is an increase in performance when using both scorers over a single scorer type, validating our earlier hypothesis. This reinforces findings from prior work (Gao et al., 2021; Kuzi et al., 2020) that hybrid methods work well. The one exception (Prime, single-field) may be challenging for single-field models, possibly due to the relatively higher number of fields in the dataset and the semantics of the fields, as we investigate more in Section 5.3. However, in the multi-field setting for Prime, we again see hybrid perform best. This provides evidence for our original motivation: that hybrid models are suitable for and positively benefits certain structured, multi-field documents.
Combining the above findings: that multi-field dense retrieval is superior, single-field BM25 can be more effective, and that hybrid works best, we experiment with mFAR 1+n, a single-field lexical, multi-field dense model. This obtains the highest average scores, including the highest score on Amazon and MAG.
5 Analysis
Next, we take a deeper look into why mFAR leads to improvements. We first verify that model is indeed adaptive to the queries by showing that query conditioning is a necessary component of mFAR. Because the field weights are naturally interpretable and controllable, we can manually set the weights to perform a post-hoc analysis of the model, which both shows us which fields of the dataset are important for the given queries and whether the model is benefiting from the dense or lexical scorers, or both, for each field. Finally, we conduct qualitative analysis to posit reasons why mFAR holds an advantage.
5.1 Is our query-conditioned adaptation necessary?
We designed mFAR with a mechanism for adaptive field selection: for a test-time query, the model makes a weighted prediction over the fields to determine which ones are important. In this section, we analyze whether this adaptation is necessary to achieve good performance. To do so, we mFAR against an ablated version which does not have the ability to predict query-specific weights but can still predict global, field-specific weights by directly learning $w_{f}^{m}$ from Equation 3. This allows the model to still emphasize (or de-emphasize) certain fields globally if they are deemed important (or unimportant).
Table 3: The test scores of mFAR All without query conditioning (QC) and the % lost without it.
Amazon MAG Prime STaRK Avg.
H@1 R@20 MRR H@1 R@20 MRR H@1 R@20 MRR H@1 R@20 MRR
mFAR All 0.412 0.585 0.542 0.490 0.717 0.582 0.409 0.683 0.512 0.437 0.662 0.545
No QC 0.346 0.547 0.473 0.428 0.662 0.528 0.241 0.596 0.368 0.338 0.602 0.456
Loss (%) -16.0 -6.5 -12.7 -12.7 -7.7 -9.3 -41.1 -12.7 -28.1 -22.6 -9.1 -16.3
In Table 3, we present the details for mFAR All and find that query conditioning is indeed necessary for performance gains across all datasets. Omitting it results in substantial losses on the metrics on each dataset and for the STaRK average. This extends to the other models too. We also find lower scores on STaRK average across the 3 metrics (H@1, R@20, MRR): -10%, -6%, -8% for mFAR Dense and -17%, -13%, -14% for the mFAR Lexical.
5.2 Which fields and scorers are important?
The interpretable design of our mFAR framework enables us to easily control the used fields and scorers after a model has been trained. Specifically, we can mask (zero) out any subset of the weights $w_{f}^{m}$ used to compute $s(q,d)$ (Equation 4). For example, setting $w_{f}^{\text{lexical}}=0$ for each $f$ would force the model to only use the dense scores for each field. We can interpret a drop in performance as a direct result of excluding certain fields or scorers, and thus we can measure their contribution (or lack thereof). In this deep-dive analysis, we study mFAR All and re-evaluate the model’s performance on each dataset after masking out entire scoring methods (lexical or dense), specific fields (title, abstract, etc), and even specific field and scoring method (e.g. title with dense scorer).
Scorers
We present results on the three STaRK datasets in Table 4. We see the performance of mFAR All on Amazon is heavily reliant on the dense scores. Knowing the results from Table 2, this may be unsurprising because mFAR Lexical did perform the worst. While the model leans similarly towards dense scores for Prime, on MAG, it relies more on the lexical scores. This shows that each dataset may benefit from a different scorer. Further, this may not be expected a priori: we would have expected Prime to benefit most from the lexical scores, as that biomedical dataset contains many initialisms and IDs that are not clearly semantically meaningful. This demonstrates the flexibility and adaptivity of mFAR to multiple scoring strategies.
From Table 2, we observe that mFAR All outperforms mFAR Dense by a small margin (0.435 vs. 0.411 for average H@$1$), and so one may suspect mFAR All is heavily relying on the dense scores. However, mFAR All with $w_{f}^{\text{lexical}}$ masked out performs substantially worse on each dataset (Table 4; 0.326 average) than mFAR Dense, suggesting that a nontrivial amount of the performance on mFAR All is attributable to lexical scores. Thus, unlike late-stage reranking or routing models for retrieval, the coexistence of dense and lexical scorers (or even individual fields) during training likely influences what the model and encoder learns.
Table 4: Performance of mFAR All with entire scoring methods masked out at test-time.
Amazon MAG Prime
Masking H@1 R@20 MRR H@1 R@20 MRR H@1 R@20 MRR
None 0.412 0.586 0.542 0.490 0.717 0.582 0.409 0.683 0.512
Dense only: $w_{f}^{\text{lexical}}=0$ 0.389 0.553 0.512 0.257 0.481 0.355 0.331 0.635 0.352
Lexical only: $w_{f}^{\text{dense}}=0$ 0.271 0.452 0.386 0.352 0.602 0.446 0.267 0.500 0.442
Fields
By performing similar analysis at a fine-grained field-level, we can identify which parts of the document are asked about or useful. For each field $f_{i}$, we can set $w_{f_{i}}^{\text{lexical}}=0$, $w_{f_{i}}^{\text{dense}}=0$, or both. We collect a few interesting fields from each dataset in Table 5, with all fields in Appendix D.
We find that behaviors vary depending on the field. For some fields (MAG’s authors, Amazon’s title), masking out one of the scorers results in almost no change. However, masking out the other one results in a sizeable drop of similar magnitude to masking out both scorers for that field. In this case, one interpretation is that $s_{\text{author}}^{\text{dense}}(q,d)$ and $s_{\text{title}}^{\text{lexical}}(q,d)$ are not useful within mFAR All.
To simplify the model, one may suggest removing any $s_{f}^{m}(q,d)$ where setting $w_{f}^{m}=0$ results in no drop. However, we cannot do this without hurting the model. In other words, low deltas do not imply low importance. For some fields (e.g. Amazon’s qa, MAG’s title, or Prime’s phenotype absent), when the lexical or dense scorers are zeroed out individually, the scores are largely unaffected. However, completely removing the field by zeroing both types of scorers results in a noticeable drop. In many cases, we observe that masking out entire fields yields a larger drop than masking out either one individually. This type of behavior could be a result of mFAR redundantly obtaining the same similarity information using different scorers. On the contrary, there is also information overlap across fields, and so in some cases, it is possible in some cases to remove entire fields, especially in Prime (e.g. enzyme) and Amazon, without substantial drops.
Table 5: For each dataset, the absolute change (delta) of masking out certain fields and scorers from mFAR All for H@$1$ and R@$20$. For each field, we zero out either the lexical scorer, the dense scorer, or both. The raw scores on all metrics for all fields in each dataset are in Appendix D.
$w_{f}^{\text{lexical}}=0$ $w_{f}^{\text{dense}}=0$ Both
Field H@1 R@20 H@1 R@20 H@1 R@20
Amazon qa 0 0 0 0 -0.031 -0.041
title 0.002 -0.003 -0.022 -0.031 -0.023 -0.024
authors -0.152 -0.117 0 0 -0.101 -0.086
MAG title -0.011 -0.003 -0.017 -0.014 -0.076 0.063
Prime phenotype absent -0.001 -0.002 0 0 -0.033 -0.030
enzyme 0 0 0 0 -0.004 -0.006
Query: Which gene or protein is not expressed in female gonadal tissue?
mFAR 2: mFAR All:
name: NUDT19P5
type: gene/protein
expression present: {anatomy: female gonad }
name: HSP90AB3P
type: gene/protein
expression absent: {anatomy: [cerebellum, female gonad]}
Query: Does Arxiv have any research papers from Eckerd College on the neutron scattering of 6He in Neutron physics?
mFAR Lexical: mFAR Dense: mFAR All:
Abstract: Abstract A new pinhole small-angle neutron scattering (SANS) spectrometer, installed at the cold neutron source of the 20 MW China Mianyang Research Reactor (CMRR) in the Institute of Nuclear Physics
Authors: Mei Peng (China Academy of Engineering Physics), Guanyun Yan (China Academy of Engineering Physics), Qiang Tian (China Academy of Engineering Physics), …
Abstract: Abstract Measurements of neutron elastic and inelastic scattering cross sections from 54Fe were performed for nine incident neutron energies between 2 and 6 MeV …
Cited Papers: Neutron scattering differential cross sections for 23 Na from 1.5 to 4.5 MeV, Neutron inelastic scattering on 54Fe
Area of Study: [Elastic scattering, Physics, Inelastic scattering, Neutron, Direct coupling, Atomic physics, Scattering, …]
Abstract: …scattering of 6He from a proton target using a microscopic folding optical potential, in which the 6He nucleus is described in terms of a 4He-core with two additional neutrons in the valence p-shell. In contrast to the previous work of that nature, all contributions from the interaction of the valence neutrons
Authors: P. Weppner (Eckerd College), A. Orazbayev (Ohio University), Ch. Elster (Ohio University)
Area of Study: [elastic scattering, physics, neutron, …, atomic physics, scattering]
Figure 3: Snippets from the highest-scoring document selected by various mFAR. Top: a single-field hybrid model (mFAR 2) vs. mFAR All. mFAR All picks correctly while mFAR 2 is possibly confused by negation in the query. Bottom: Snippets from configurations of mFAR with access to different scorers. Only mFAR All correctly makes use of both lexical and semantic matching across fields.
5.3 Qualitative Analysis
Multi-field gives semantic meaning for a choice of field, as compared to single-field.
In Figure 3 (top), the query is looking for either a gene or protein that is not expressed. With mFAR All, the retriever matches a longer text more accurately than the single-field hybrid retriever doesours Hybridand mFAR 2 correctly match female gonad. However, mFAR selects the field that refers to the absence of an expression, which the model learns. In the mFAR 2, because the lexical scorer cannot distinguish between present and absent and the document embedding also failed to do so, it incorrectly ranks a negative document higher.
Hybrid excels when both lexical matching and semantic similarity is required.
In Figure 3 (bottom), mFAR All has the advantage over mFAR Dense by having the ability to lexically match Eckerd College. Furthermore, mFAR All is still able to semantically match the abstract of the document. While mFAR Dense also finds a close fit, it is unable to distinguish this incorrect but similar example from the correct one.
We likewise observe the drawbacks of a lexical-only scoring. One limitation of BM25 is that the frequency of successive term matching results in increased scores. Because Physics is a keyword with high frequency in the authors list, it results in a high score for this document even though it is not used in the same sense semantically. On the other hand, mFAR All correctly matches the specific institution because the final scores are based on a weighted combination of lexical and dense scorers, which may reduce or the impact of high lexical scores.
6 Related Work
Multi-field and Hybrid Retrieval.
Multi-field retrieval has previously been explored in the context of sparse methods (Robertson et al., 2004; Zaragoza et al., 2004) which can involve extensions of the lexical-based BM25 (Robertson et al., 1994), learned sparse representations (Zamani et al., 2018), or Bayesian approaches (Piwowarski & Gallinari, 2003). Recently, Lin et al. (2023) approached a similar task with a primary focus on query decomposition and rewriting. They also propose a weighted combination of “expert” retrievers, though their weights are hand-picked and each (dense) retriever is trained independently. In mFAR, we finetune the entire system end-to-end with a single shared encoder and our weights are learned, which allows us to scale to dozens of fields.
The combination of both types of lexical and dense scorers has previously been found to be complementary, leading to performance gains (Gao et al., 2021; Kuzi et al., 2020). Notably, Kuzi et al. (2020) points out that long documents are challenging for lexical-based methods and suggests document chunking as a possible remedy in future work. We implicitly segment the document by taking advantage of the multi-field structure inherently present in documents, and unlike those past works, our work is the first to demonstrate the strength of hybrid-based methods in a multi-field setting.
While we focus on documents with multi-field structure, prior work has focused on documents with other types of structure. For example, Chen et al. (2020b) design a model for table search, while Husain et al. (2019) propose the task for code search. Reddy et al. (2022) introduce an Amazon shopping dataset with short search queries and proposes an embeddings-based method for document ranking, relevance classification, and product substitutions. We already use an Amazon dataset within STaRK, include several embeddings-based baselines, and are more interested in queries that relate to multiple fields, which is not the focus of their task. The dataset from Lin et al. (2023) is multimodal; our work is unimodal: we focus first on natural language and next the question of lexical and dense scorers.
Some prior works have incorporated document structure as part of pretraining. Su et al. (2023) generates pseudo-queries based on Wikipedia formatting, while Li et al. (2023) introduces an alignment objective to increase similarity between the structured part of the dataset and its corresponding natural-language description. Both of these pretraining-focused methods assume some redundancy in the underlying data, which STaRK, and multi-field documents in general, do not guarantee, and we focus on introducing a post-training method that does not modify the underlying model pretraining scheme.
7 Conclusion
We present mFAR, a novel framework for retrieval over multi-field data by using multiple scorers, each independently scoring the query against a part (field) of a structured document. These scorers can be lexical-based or dense-based, and each field can be scored by both types. We introduce an interpretable and controllable query-conditioned predictor of weights used to adaptively sum over these scores. On three large-scale datasets, we find that mFAR can achieve significant performance gains over existing methods due to multi-field advantages and the inclusion of a hybrid combination of scorers, leading to state-of-the-art performance. Through our analysis, we find that the best models benefit from both access to both scorers and the ability to weight each field conditioned on the query, further verifying our method.
Our primary goal is to study the challenging and emerging problem of retrieval for multi-field structured data and to introduce a flexible framework to approach it. Having laid the groundwork, future work can include more specialized individual scorers, scale up to more scorers in other modalities like vision or audio, and add other algorithmic improvements to the weighted integration of scores across scorers. Then, mFAR would be a step towards retrieval of any type of content, which can further aid applications for general search or agent-based RAG.
Appendix A Dataset
A.1 Preprocessing
Technically, STaRK is a dataset of queries over knowledge graphs. The baselines, Wu et al. (2024b), create a linearized document for each node, which omits some edge and multi-hop information that is available in the knowledge graph. AvaTaR (Wu et al., 2024a) operates directly on the knowledge graph. As we want to operate over structured documents, we need a preprocessing step either on the linearized documents or by processing the graph.
Because parsing documents is error-prone, we decide to reproduce the document creation process from (Wu et al., 2024b). We start with all of the original dataset from STaRK, which come in the form of queries and the associated answer ids in the knowledge graph. Each query requires a combination of entity information and relation information from their dataset to answer. However, each dataset handles the entity types differently. The answer to every query for Amazon is the product entity. For MAG, the answer is the paper entity. However, Prime has a list of ten possible entities that can answer the query, so we include all ten as documents.
We create our set of documents based on the directional paths taken in their knowledge graph; if there are more than single hop relations, then we take at most two hops for additional entities and relations. For Amazon, since the queries are at most one hop, we do not include additional node information. MAG and Prime, however, can include more queries with more than two hops, so we include information about additional relations and nodes for each document in our dataset.
A.2 Fields
We include the list of fields that we used in this work in Table 6. Not every single field available in the STaRK knowledge graph (Wu et al., 2024b) is used because some are not used in the baseline and we try to match the baselines as closely as possible. We make some cosmetic changes for space and clarity in the examples in the main body of this paper, including uppercasing field names and replacing underscore with spaces. We also shorten “author___affiliated_with___institution”, “paper___cites___paper” to “Papers Cited”, “paper___has_topic___field_of_study” to “Area of Study” and expand “type” to “Entity Type.”
Table 6 also lists some information about the length distribution of each field, as measured by the Contriever tokenizer. This is useful to know how much information might be lost to the limited window size of Contriever. Furthermore, we list the maximum sequence length used by the dense scorer of mFAR both during training and at test-time. The trade off for sequence length is batch size with respect to GPU memory usage. Our lexical baseline (BM25) does not perform any truncation.
Table 6: The datasets and list of fields, $\mathcal{F}$, used in this work, along with basic length statistics of the content of those fields. The length of the $k$th-%ile longest value is listed. For example, $k=50$ would be the median length. The MSL is the maximum sequence length threshold we chose for that field in mFAR based on either the maximum window size of the encoder (512) or based on covering most ($>99\%$) documents within the corpus..
Length in Contriever tokens ($k^{\text{th}}$%-ile)
Dataset Field 90 95 99 99.9 Max MSL
Amazon also_buy 64 217 908 3864 50176 512
also_view 86 189 557 1808 21888 512
brand 7 8 10 12 35 16
description 207 289 446 1020 5038 512
feature 130 171 305 566 1587 512
qa 4 4 5 698 1873 512
review 1123 2593 12066 58946 630546 512
title 28 34 48 75 918 128
MAG abstract 354 410 546 775 2329 512
author___affiliated_with___institution 90 121 341 18908 46791 512
paper___cites___paper 581 863 1785 4412 79414 512
paper___has_topic___field_of_study 49 52 57 63 90 64
title 31 34 44 62 9934 64
Prime associated with 10 35 173 706 4985 256
carrier 3 4 4 13 2140 8
contraindication 4 4 66 586 3481 128
details 329 823 2446 5005 12319 512
enzyme 4 4 12 63 5318 64
expression absent 4 8 29 77 12196 64
expression present 204 510 670 18306 81931 512
indication 4 4 25 146 1202 32
interacts with 93 169 446 1324 55110 512
linked to 3 4 4 57 544 8
name 17 21 38 74 133 64
off-label use 3 4 4 56 727 8
parent-child 49 70 168 714 18585 256
phenotype absent 3 4 4 33 1057 8
phenotype present 20 82 372 1931 28920 512
ppi 36 125 438 1563 22432 512
side effect 4 4 93 968 5279 128
source 5 6 6 7 8 8
synergistic interaction 4 4 4800 9495 13570 512
target 4 9 33 312 5852 64
transporter 3 4 4 41 2721 8
type 7 8 8 9 9 8
Appendix B Implementation Details
During training, we sample 1 negative example per document. Along with in-batch negatives, this results in $12b-$ 1 negative samples for a batch size of $b$. This negative document is sampled using Pyserini Lucene [^4] https://github.com/castorini/pyserini: 100 nearest documents are retrieved, of which the postive documents are removed. The top negative document is then sampled among that remaining set. We apply early stopping on validation loss with a patience of 5. We set $\tau$ = 0.05 and train with DDP on 8x NVIDIA A100s. Contriever is a 110M parameter model, and the additional parameters added through $G$ is negligible ($768|\mathcal{F}|$), scaling linearly in the number of fields.
We use separate learning rates (LRs) for finetuning the encoder and for the other parameters. Specifically, we searched over learning rates [5e-6, 1e-5, 5e-5, 1e-4] for the encoder and [1e-3, 5e-3, 1e-2, 5e-2, 1e-1] for the parameters in $G(q,f,m)$ which consist of $\mathbf{a}_{f}^{m}$ and $\gamma_{f}^{m}$ and $\beta_{f}^{m}$ from batch normalization. The main grid search was conducted over the bolded values, although we found 5e-3 to be effective for $G(q,f,m)$ for Amazon. We otherwise follow the default settings for both the optimizer (AdamW, dropout, etc. from sentence-transformers 2.2.2) and batch normalization (PyTorch 2.4.0). As mentioned, whether to apply batch normalization at all was also a hyperparameter searched over: we found it useful in the hybrid setting. Our implementation uses Pytorch Lightning [^5] https://lightning.ai/ and sentence-transformers (Reimers & Gurevych, 2019). We use a fast, python-based implementation of BM25 as our lexical scorer (Lù, 2024). [^6] https://github.com/xhluca/bm25s The best hyperparameters for each of our models in this work are listed in Table 7. In the case where there is only a single field (last two sections), the adaptive query conditioning is not needed.
At inference, we retrieve the top-$100$ results per field to form a candidate set, and we compute the full scores over this candidate set to obtain our final ranking.
Table 7: The hyperparameters used for each of the runs in this work.
Model Dataset Referenced in Encoder LR $G()$ LR Batch norm?
mFAR All Amazon Table 2, 3, most tables/figures 1e-5 5e-3 no
MAG 5e-5 1e-2 yes
Prime 5e-5 1e-2 yes
mFAR Dense Amazon Table 2, Figure 3 1e-5 5e-3 no
MAG 5e-5 5e-2 no
Prime 1e-5 1e-2 no
mFAR Lexical Amazon Table 2, Figure 3 1e-5 5e-3 yes
MAG 1e-5 1e-2 yes
Prime 5e-5 1e-1 yes
mFAR 2 Amazon Table 2, Figure 3 1e-5 1e-2 no
MAG 5e-5 5e-3 yes
Prime 5e-5 5e-3 yes
mFAR 1+n Amazon Table 2 1e-5 1e-2 no
MAG 5e-5 5e-3 yes
Prime 1e-5 5e-3 yes
Contriever-FT Amazon Table 2 5e-5 n/a n/a
MAG 1e-5 n/a n/a
Prime 5e-5 n/a n/a
Appendix C Full Results on STaRK
We present comprehensive results on the test split of Amazon, MAG, and Prime.
Full test results and comparison
Here, we report the same results as in the main section, but we also include H@$5$, to be exhaustive STaRK, in addition to our existing metrics. In Table 8, we show the test results with the included additional metric. In Table 9, we also include the average as a separate table. Here, we find that mFAR still does on average better than the other baselines on the structured datasets, even against the strong lexical BM25 baseline.
Table 8: Similar to Table 2, we instead include H@5 and show the average as a separate table, over the test split. We include the same baselines and generally find that H@5 also follows the same trends. The average over all datasets can be seen in Table 9. * In Wu et al. (2024b), these are only run on a random 10% subset.
Amazon MAG Prime
Model H@1 H@5 R@20 MRR H@1 H@5 R@20 MRR H@1 H@5 R@20 MRR
ada-002 0.392 0.627 0.533 0.542 0.291 0.496 0.484 0.386 0.126 0.315 0.360 0.214
multi-ada-002 0.401 0.650 0.551 0.516 0.259 0.504 0.508 0.369 0.151 0.336 0.381 0.235
Claude3* 0.455 0.711 0.538 0.559 0.365 0.532 0.484 0.442 0.178 0.369 0.356 0.263
GPT4* 0.448 0.712 0.554 0.557 0.409 0.582 0.486 0.490 0.183 0.373 0.341 0.266
BM25 0.483 0.721 0.584 0.589 0.471 0.693 0.689 0.572 0.167 0.355 0.410 0.255
Contriever-FT 0.383 0.639 0.530 0.497 0.371 0.594 0.578 0.475 0.325 0.548 0.600 0.427
AvaTaR (agent) 0.499 0.692 0.606 0.587 0.444 0.567 0.506 0.512 0.184 0.367 0.393 0.267
mFAR Lexical 0.332 0.569 0.491 0.443 0.429 0.634 0.657 0.522 0.257 0.455 0.500 0.347
mFAR Dense 0.390 0.659 0.555 0.512 0.467 0.678 0.669 0.564 0.375 0.620 0.698 0.485
mFAR 2 0.574 0.814 0.663 0.681 0.503 0.717 0.721 0.603 0.227 0.439 0.495 0.327
mFAR All 0.412 0.700 0.585 0.542 0.490 0.696 0.717 0.582 0.409 0.628 0.683 0.512
mFAR 1+n 0.565 0.816 0.659 0.674 0.511 0.731 0.748 0.611 0.359 0.603 0.650 0.469
Table 9: The averages for Table 8.
Averages
Model H@1 H@5 R@20 MRR
ada-002 0.270 0.479 0.459 0.381
multi-ada-002 0.270 0.497 0.480 0.373
Claude3* 0.333 0.537 0.459 0.421
GPT4* 0.347 0.556 0.460 0.465
BM25 0.374 0.590 0.561 0.462
Contriever-FT 0.360 0.594 0.569 0.467
AvaTaR (agent) 0.376 0.542 0.502 0.455
mFAR Lexical 0.339 0.553 0.549 0.437
mFAR Dense 0.411 0.652 0.641 0.520
mFAR 2 0.435 0.656 0.626 0.537
mFAR All 0.437 0.675 0.662 0.545
mFAR 1+n 0.478 0.717 0.686 0.585
Appendix D Full results for field masking
We include full scores for masking each field and scorer for Amazon in Table 10, MAG in Table 11, and Prime in Table 12. The first row “—” is mFAR All without any masking and repeated three times as a reference. The final row “all” is the result of masking out all the lexical scores (or all the dense scores). It does not make sense to mask out all scores, as that would result in no scorer.
Based on our findings in Table 11, all fields in MAG are generally useful, as all instances of zeroing out the respect fields results in a performance drop. Despite this finding with MAG, not all fields are as obviously important in other datasets. For Table 12, Prime has a notable number of fields that do not contribute to the final ranking when both scorers are masked out. And for Amazon, in Table 10, we surprisingly find that fields like “description” and “brand” have little effect. This is a reflection on both the dataset (and any redundancies contained within) and on the distribution of queries and what they ask about.
Table 10: Test scores on Amazon after masking out each field and scorer of the mFAR All at test-time.
Amazon $w_{f}^{\text{lexical}}=0$ $w_{f}^{\text{dense}}=0$ Both
Masked field H@1 H@5 R@20 MRR H@1 H@5 R@20 MRR H@1 H@5 R@20 MRR
0.412 0.700 0.586 0.542 0.412 0.700 0.586 0.542 0.412 0.700 0.586 0.542
also_buy 0.407 0.690 0.578 0.534 0.410 0.696 0.586 0.540 0.403 0.678 0.578 0.530
also_view 0.420 0.695 0.576 0.542 0.414 0.696 0.581 0.542 0.395 0.677 0.565 0.522
brand 0.410 0.699 0.585 0.540 0.397 0.692 0.575 0.528 0.400 0.686 0.570 0.526
description 0.417 0.699 0.587 0.542 0.410 0.692 0.580 0.540 0.413 0.680 0.576 0.535
feature 0.412 0.700 0.581 0.537 0.398 0.680 0.570 0.524 0.410 0.680 0.562 0.531
qa 0.412 0.700 0.586 0.542 0.412 0.700 0.586 0.542 0.381 0.636 0.545 0.499
review 0.410 0.696 0.583 0.541 0.398 0.680 0.575 0.526 0.384 0.666 0.548 0.510
title 0.414 0.685 0.583 0.535 0.390 0.650 0.555 0.508 0.389 0.672 0.562 0.516
all 0.389 0.660 0.553 0.512 0.271 0.518 0.452 0.386
Table 11: Test scores on MAG after masking out each field and scorer of the mFAR All at test-time. Due to space, we truncate some field names, refer to Table 6 for the full names.
MAG $w_{f}^{\text{lexical}}=0$ $w_{f}^{\text{dense}}=0$ Both
Masked field H@1 H@5 R@20 MRR H@1 H@5 R@20 MRR H@1 H@5 R@20 MRR
0.490 0.696 0.717 0.582 0.490 0.696 0.717 0.582 0.490 0.696 0.717 0.582
abstract 0.469 0.681 0.707 0.565 0.393 0.616 0.651 0.494 0.430 0.636 0.659 0.526
author affil… 0.338 0.555 0.600 0.439 0.490 0.696 0.717 0.582 0.389 0.595 0.631 0.485
paper cites… 0.458 0.660 0.655 0.551 0.484 0.685 0.708 0.576 0.424 0.650 0.668 0.526
paper topic… 0.459 0.671 0.695 0.554 0.491 0.695 0.717 0.582 0.398 0.617 0.650 0.499
title 0.479 0.686 0.714 0.573 0.473 0.676 0.703 0.565 0.414 0.633 0.654 0.513
all 0.257 0.462 0.481 0.355 0.352 0.561 0.602 0.446
Table 12: Test scores on Prime after masking out each field and scorer of the mFAR All at test-time. Due to space, we shorten some field names, refer to Table 6 for the full names.
Prime $w_{f}^{\text{lexical}}=0$ $w_{f}^{\text{dense}}=0$ Both
Masked field H@1 H@5 R@20 MRR H@1 H@5 R@20 MRR H@1 H@5 R@20 MRR
0.409 0.627 0.683 0.512 0.409 0.627 0.683 0.512 0.409 0.627 0.683 0.512
associated with 0.392 0.610 0.670 0.495 0.407 0.624 0.680 0.510 0.399 0.618 0.672 0.502
carrier 0.409 0.627 0.683 0.512 0.409 0.627 0.683 0.512 0.403 0.621 0.678 0.506
contraindication 0.409 0.627 0.683 0.512 0.409 0.627 0.683 0.512 0.380 0.587 0.652 0.479
details 0.386 0.606 0.670 0.488 0.363 0.569 0.619 0.458 0.388 0.601 0.661 0.489
enzyme 0.409 0.627 0.683 0.512 0.409 0.627 0.683 0.512 0.405 0.623 0.677 0.508
expression abs. 0.408 0.627 0.683 0.511 0.392 0.607 0.664 0.494 0.403 0.622 0.678 0.506
expression pres. 0.409 0.627 0.683 0.512 0.409 0.627 0.683 0.512 0.400 0.617 0.675 0.502
indication 0.407 0.627 0.682 0.511 0.398 0.613 0.663 0.498 0.392 0.611 0.665 0.495
interacts with 0.403 0.624 0.681 0.507 0.406 0.626 0.682 0.510 0.403 0.622 0.674 0.506
linked to 0.409 0.627 0.683 0.512 0.409 0.627 0.683 0.512 0.383 0.601 0.661 0.486
name 0.410 0.628 0.684 0.513 0.407 0.627 0.681 0.510 0.407 0.622 0.674 0.507
off-label use 0.409 0.627 0.683 0.512 0.409 0.627 0.683 0.512 0.379 0.602 0.662 0.482
parent-child 0.385 0.619 0.680 0.494 0.391 0.613 0.663 0.495 0.386 0.601 0.663 0.487
phenotype abs. 0.408 0.625 0.681 0.511 0.409 0.627 0.683 0.512 0.376 0.591 0.653 0.477
phenotype pres. 0.405 0.619 0.675 0.506 0.409 0.627 0.683 0.512 0.393 0.609 0.669 0.495
ppi 0.403 0.622 0.678 0.506 0.409 0.627 0.683 0.512 0.399 0.617 0.671 0.502
side effect 0.409 0.627 0.683 0.512 0.409 0.627 0.683 0.512 0.405 0.624 0.680 0.508
source 0.409 0.627 0.683 0.512 0.409 0.627 0.683 0.512 0.397 0.614 0.671 0.499
synergistic int. 0.408 0.627 0.682 0.511 0.409 0.627 0.683 0.512 0.381 0.597 0.659 0.483
target 0.407 0.627 0.683 0.511 0.394 0.613 0.662 0.497 0.397 0.617 0.671 0.501
transporter 0.409 0.627 0.683 0.512 0.409 0.627 0.683 0.512 0.406 0.624 0.679 0.509
type 0.409 0.627 0.683 0.512 0.403 0.625 0.681 0.507 0.396 0.615 0.669 0.498
all 0.342 0.554 0.624 0.442 0.267 0.450 0.500 0.352